How to calculate the prevalence ratio (vs. 7 hours) with this table: Exhibit I.2 Age < 60 Age > 60 "# Cases" "# Risk" Rate/100 "PR (vs.7 hours" "# Cases" "# Risk" Rate/100 "PR (vs.7 hours" ≤ 5 hours 147 1,911 7.69 157 625 25.12 6 hours 160 5,000 3.2 258 1,369 18.85 7 hours 119 7,154 1.66 315 1,969 16 8 hours 145 6,914 2.1 494 2,774 17.81 ≥ 9 hours 57 1,526 3.74 294 1,155 25.45
Answer #1

Prevalence rate is the total ( new and pre-existing ) number of cases of a disease that are present in a particular population at a given time.

Formula = Number of old and new cases of a disease / Total population at risk of getting that disease × 1000

For population less than 60years of age

Time less than or equal to 5 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 147/ 1911 × 1000

= 7.69

Time is equal to 6 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 160/ 5000×1000

= 3.2

Time is equal to 7 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 119/ 7154

= 166

Time = 8 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 145/6914

= 2.1

Time more than or equal to 9 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 57/1526×1000

= 3.74

For population more than 60years of age

Time less than or equal to 5hours sleep ,prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 157/ 625×1000

= 25.12

Time is equal to 6 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 258/ 1369× 1000

= 18.85

Time is equal to 7 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 315/1969×1000

= 16

Time = 8 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 494/ 2774× 1000

= 17.81

Time is more than or equal to 9 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 294/ 1155 × 1000

= 25.45